OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (amistre64):

parallel lines differ at most by a constant; all of these lines have different constants at the end of them

OpenStudy (amistre64):

it might be simplest to plug in the value of the point and see which one it fits

OpenStudy (amistre64):

unless you interact, ill assume that makes perfect sense to you .....

OpenStudy (anonymous):

Not really

OpenStudy (amistre64):

do you know how to read a point?

OpenStudy (amistre64):

points are ordered pairs, in this case denoted as (x,y) as such the point (6,-2) tells us that we want x to be 6 and y to be -2 substitute those values into the equations for x and y

OpenStudy (amistre64):

lets take a look at the first option: y = x + 8 plugging in the point values we get: -2 = 6 + 8 -2 = 14 ; which is a false statement so it cant be this one

OpenStudy (amistre64):

can you show me how the second option would work out?

OpenStudy (anonymous):

do you divid -2 on both side of the equal?

OpenStudy (amistre64):

no. this problem only requires basic adding and subtracting skills

OpenStudy (amistre64):

y = x - 4 ; just plug in the point values (6,-2) -2 = 6 - 4 ; and work the math -2 = 2 ; the math ends up a false statement since -2 does not equal 2 option 2 is not it